OpenStudy (anonymous):

Which of the following statements have the same result? Explain each step in solving each one. I.f(2) when f(x) = 3x + 2 II.f-1(3) when f(x) = 2 x minus 7, all over 3 III.2y + 14 = 4y - 2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I really could use the help, I need a good grade:/

ganeshie8 (ganeshie8):

solve the variable in each statement

OpenStudy (anonymous):

I don't get how to do that..that's what I need the help on

ganeshie8 (ganeshie8):

I.f(2) when f(x) = 3x + 2 find f(2) by putting x = 2 f(x) = 3x+2 f(2) = ?

OpenStudy (anonymous):

f(2)= 2?

OpenStudy (anonymous):

f of x is 3x+2 then f of 2 is 3(2)+2=8

OpenStudy (anonymous):

Okay so f(2) = 3(2) + 2 = 8 now what?

OpenStudy (loser66):

so you have 1/ f(2) = 8 let it there now, 2/ \(f^{-1}(3) \) when \(\large f(x )= \frac{2x-7}{3}\) you have to find \(f^{-1}(x) \) which is \(\large \frac{3x+7}{2}\) and then \(f^{-1} (3) = 8\) So, 1/ = 2 2/ = 8 now, 3/ solve for y : 2y+14 = 4y -2 16 = 2y y = 8 Summary : 1/ = 2 2/ = 8 3/ =8 conclusion: 2/ and 3/ have the same solution. Read it carefully, hopefully you can follow and know how to work on this kind of problem from now on. Good luck

OpenStudy (loser66):

oh, mistake, f(2 ) =8 , they are all equal. I am sorry.

OpenStudy (anonymous):

So all the soutions are equal and have the same result?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Thanks, so much

OpenStudy (anonymous):

@Loser66 is wrong. Only 1 and 3 have the same result. I. 3(2)+2 6+2 8 II. 2(3)-7/3 6-7/3 -1/3 -3 III. 2y+14 = 4y-2 16 = 2y 8 = y ANSWER: I and III have the same result.

OpenStudy (loser66):

@Nico_dangond II is \(f^{-1}(3) \) , not just f (3) so that ,you must find inverse of f (x) first.